# Controlling Space and Time with Diffusion Models

### v1

@@ Appendix A Calibrated RealEstate10K
Appendix B Neural Architecture
@@
As hinted by Figure 2, 4DiM uses a UViT architecture following Hoogeboom et al. [2023]. Compared to the commonly employed UNet architecture [Ronneberger et al., 2015], UViT uses a transformer backbone at the bottleneck resolution with no convolutions, leading to improved accelerator utilization. Information across frames mixes exclusively in temporal attention blocks, following [Ho et al., 2022b]. We find that this is key to keep computational costs within reasonable limits compared to, say, 3D attention [Shi et al., 2023]. Because the temporal attention blocks have a limited sequence length (32 frames), we insert them at all UViT resolutions. To limit memory usage, we use per-frame self-attention blocks only at the 16x16 bottleneck resolution where the sequence length is the product of the current resolution’s dimensions.
@@
Our transformer blocks combine helpful choices from several prior work. We use parallel attention and MLP blocks following Dehghani et al. [2023], as we found in our early experiments that this leads to slightly better hardware utilization while achieving almost identical sample quality. We also inject conditioning information in a similar fashion to Peebles and Xie [2023], though we use fewer conditioning blocks due to the use of the parallel attention + MLP. We additionally emply query-key normalization following the findings of Gilmer et al. for improved training stability, but with RMSNorm [Zhang and Sennrich, 2019] for simplicity.
@@
We use the same positional encodings for diffusion noise levels and relative timestamps as [Saharia et al., 2022b]. For relative poses, we follow 3DiM [Watson et al., 2022], and condition generation with per-pixel ray origins and directions, as originally proposed by SRT [Sajjadi et al., 2022] in a regressive setting. Xiong et al. [2023] has compared this choice to conditioning via encoded extrinsics and focal lengths a-la Zero-1-to-3 [Liu et al., 2023b], finding it advantageous. While a thorough study of different encodings is missing in the literature, we hypothesize that rays are a natural choice as they encode camera intrinsics in a way that is independent of the target resolution, yet gives the network precise, pixel-level information about what contents of the scene are visible and which are outside the field of view and therefore require extrapolation by the model. Unlike Plucker coordinates, rays additionally preserve camera positions which is key to deal with occlusions in the underlying 3D scene.
Appendix C Compute, training time, and sharding
@@
We train 8-frame 4DiM models for 1M steps. Using 64 TPU v5e chips, we achieve a throughput of approximately 1 step per second with a batch size of 128. The model has 2.6B parameters, and available HBM is maximized with various strategies: first, we use bfloat16 activations (but still use float32 weights to avoid instabilities). We also use FSDP (i.e., zero-redundancy sharding [Rajbhandari et al., 2020] with delayed and rematerialized all-gathers). We then finetune this model to its final 32-frame version. This strategy allows us to leverage large batch-size pre-training, which is not possible with too many frames because the amount of activations stored for backpropagation scales linearly respect the number of frames per training example. The model is finetuned with the same number of chips, albeit only for 50,000 steps and at batch size of 32. This allows us to shard the frames of the video and use more than one chip per example at batch size 1. When using temporal attention, we simply all-gather the keys and values and keep the queries sharded over frames (one of the key insights from Ring Attention [Liu et al., 2023a]), though we don’t decompose the computation of attention further as we find we have sufficient HBM to fully parallelize over all-gathered keys and values. FSDP is also enabled on the frame axis to maximize HBM savings so the number of shards is truly the number of chips (as opposed to the number of batch-parallel towers). This is possible because frame sharding requires an all-reduce by mean on the loss over the frame axis. Identically to zero-redundancy sharding, this can be broken down into a reduce-scatter followed by an all-gather.
Appendix D Further details on proposed metrics
@@
We follow Yu et al. [2023a] and compute TSED only for contiguous pairs in each view trajectory, and discard sequences where we find less than 10 two-view keypoint matches. We use a threshold of 2.0 in all our experiments. For our proposed SfM distances, in order to get a scale-invariant metric, we first align the camera positions predicted by COLMAP by relativizing the predicted and original poses with respect to the first conditioning frame. This should resolve rotation ambiguity. Then, we resolve scale ambiguity by analytically solving for the least squares optimization problem that aligns the original positions to re-scaled COLMAP camera positions. Finally, we report the relative error in positions under the norm, i.e., we normalize by the norm of the original camera positions.
Appendix E Further details on baselines from prior work
@@
Appendix F Samples

### v2

@@ Appendix A More details on calibrating and filtering RealEstate10K
Appendix B Pushing Pareto fronts with multi-guidance
To assess the effectiveness of multi-guidance, we conducted a two-stage evaluation. First, we evaluated samples generated with various ‘standard’ guidance weights, where all conditioning variables (image, pose, and timestamp) were treated uniformly (illustrated in blue in Figure 7). Through qualitative and quantitative analysis, we identified the optimal standard guidance weight. Next, while holding this image guidance weight constant, we systematically varied the pose or timestamp guidance weights (shown in red). This procedure was applied to two tasks: single-image 3D generation using the cRE10K dataset, and video extrapolation from the first two frames using the DAVIS dataset. As demonstrated in Figure 7, our findings indicate that for video extrapolation, a slightly stronger timestamp guidance enhances performance in both dynamics and FID. For 3D generation, a slightly stronger pose guidance improves pose alignment (measured by TSED) with minimal impact on FID.
Appendix C Further details on neural architecture
@@
As hinted by Figure 2, 4DiM uses a UViT architecture following Hoogeboom et al. (2023). Compared to the commonly employed UNet architecture (Ronneberger et al., 2015), UViT uses a transformer backbone at the bottleneck resolution with no convolutions, leading to improved accelerator utilization. Information across frames mixes exclusively in temporal attention blocks, following (Ho et al., 2022b). We find that this is key to keep computational costs within reasonable limits compared to, say, 3D attention (Shi et al., 2023). Because the temporal attention blocks have a limited sequence length (32 frames), we insert them at all UViT resolutions. To mitigate the quadratic cost of attention, we use per-frame self-attention blocks only at the 16x16 bottleneck resolution where the sequence length is the product of the current resolution’s dimensions. And to further minimize the amount of activations stored for backpropagation, we use only two UNet blocks when the resolution is at 32x or beyond, with hidden sizes (channels) decreasing in powers of two to a minimum of 256. Convolutional kernels at the 32x resolution have 3x3 kernels, and we increase the kernel size additively by 2 for higher resolutions to a maximum of 7x7. Our base model uses 16 transformer blocks at the bottleneck resolution of hidden size 128 with 16 heads, and our super-resolution model uses 8 transformer blocks of hidden size 128 with 8 heads.
@@
Our transformer blocks combine helpful choices from several prior work. We use parallel attention and MLP blocks following Dehghani et al. (2023), as we found in our early experiments that this leads to slightly better hardware utilization while achieving almost identical sample quality. We also inject conditioning information in a similar fashion to Peebles & Xie (2023), though we use fewer conditioning blocks due to the use of the parallel attention + MLP. We additionally emply query-key normalization following the findings of Gilmer et al. for improved training stability, but with RMSNorm (Zhang & Sennrich, 2019) for simplicity.
@@
We use the same positional encodings for diffusion noise levels and relative timestamps as (Saharia et al., 2022b). For relative poses, we follow 3DiM (Watson et al., 2022), and condition generation with per-pixel ray origins and directions, as originally proposed by SRT (Sajjadi et al., 2022) in a regressive setting. Xiong et al. (2023) has compared this choice to conditioning via encoded extrinsics and focal lengths a-la Zero-1-to-3 (Liu et al., 2023b), finding it advantageous. While a thorough study of different encodings is missing in the literature, we hypothesize that rays are a natural choice as they encode camera intrinsics in a way that is independent of the target resolution, yet gives the network precise, pixel-level information about what contents of the scene are visible and which are outside the field of view and therefore require the model to extrapolate. Unlike Plucker coordinates, which parametrize lines in 3D space, rays additionally preserve information about camera positions, which is key to deal with occlusions in the underlying 3D scene.
Appendix D Further details on compute, training time, and sharding
We train 8-frame 4DiM models for 1M steps. Using 64 TPU v5e chips, we achieve a throughput of approximately 1 step per second with a batch size of 128. All 8-frame models we train (i.e. base and super-resolution models) use this batch size, and we train all models with an Adam (Kingma & Ba, 2014) learning rate of 0.0001 linearly warmed up for the first 10,000 training steps, which we found was the best peak value in early sweeps. No learning rate decay is used. We follow Ho et al. (2020) and keep an exponential moving average of the parameters with a decay rate of 0.9999 to use at inference time for improved sample quality.
@@ Appendix D Further details on compute, training time, and sharding
Available HBM is maximized with various strategies: first, we use bfloat16 activations (but still use float32 weights to avoid instabilities). We also use FSDP (i.e., zero-redundancy sharding (Rajbhandari et al., 2020) with delayed and rematerialized all-gathers). We then finetune this model to its final 32-frame version. This strategy allows us to leverage large batch-size pre-training, which is not possible with too many frames because the amount of activations stored for backpropagation scales linearly with respect the number of frames per training example. The model is finetuned with the same number of chips, albeit only for 50,000 steps and at batch size of 32. This allows us to shard the frames of the video and use more than one chip per example at batch size 1. When using temporal attention, we simply all-gather the keys and values and keep the queries sharded over frames (one of the key insights from Ring Attention (Liu et al., 2023a)), though we don’t decompose the computation of attention further as we find we have sufficient HBM to fully parallelize over all-gathered keys and values. FSDP is also enabled on the frame axis to maximize HBM savings so the number of shards is truly the number of chips (as opposed to the number of batch-parallel towers). This is possible because frame sharding requires an all-reduce by mean on the loss over the frame axis. Identically to zero-redundancy sharding, this can be broken down into a reduce-scatter followed by an all-gather.
Appendix E Further details on proposed metrics
@@
We follow Yu et al. (2023a) and compute TSED only for contiguous pairs in each view trajectory, and discard sequences where we find less than 10 two-view keypoint matches. We use a threshold of 2.0 in all our experiments. For our proposed SfM distances, in order to get a scale-invariant metric, we first align the camera positions predicted by COLMAP by relativizing the predicted and original poses with respect to the first conditioning frame. This should resolve rotation ambiguity. Then, we resolve scale ambiguity by analytically solving for the least squares optimization problem that aligns the original positions to re-scaled COLMAP camera positions. Finally, we report the relative error in positions under the norm, i.e., we normalize by the norm of the original camera positions.
Appendix F Further details on baselines from prior work
@@
For ZeroNVS, we use its diffusion model to evaluate novel view synthesis on RE10K and LLFF. Both input and output images are . We use default hyper-parameters as in the open sourced configuration, except the ‘scene scale’ parameter used for dealing with length scale ambiguity. We sweep over scale in and choose for RE10K and for LLFF.
Appendix G Qualitative comparison of 360∘ generation vs ZeroNVS
We employ the ZeroNVS pipeline with score distillation sampling (SDS) to generate 360∘ views conditioned on a single image. For each scene, we use the default scene-generation configuration as open sourced by ZeroNVS, as adjusting camera_distance and elevation_angle yield worse and incoherent generation. We adjust the scale parameter for each scene, and use for the three scenes in Figure 8. Due to the computationally intensive nature of 360∘ generation with ZeroNVS (2-3 hours per scene), exhaustive hyperparameter tuning is impractical. We also observe another limitation: volume rendering will fail if the camera positions are not allowed to vary (or even if camera_distance is too small, e.g., 0.5), i.e., it is not feasible to keep the camera position fixed and only employ rotation. This occurs because there is no parallax and it is therefore not possible to infer depth/density (rotation alone only warps pixels). Figure 8 presents a comparison between 4DiM and ZeroNVS outputs, showing the superior sharpness and richer content achieved by 4DiM.
Appendix H Comparisons of 4DiM vs prior frame interpolation methods
We compare 4DiM against prior works on the task of frame interpolation (generating a video with large amounts of motion given the first and last frames). We use the Davis dataset as the benchmark and consider a generative baseline, LDMVFI (Danier et al., 2024), as well as the well-known (but non-generative) works FILM (Reda et al., 2022) and RIFE (Huang et al., 2022).
Below, we report FID scores of the middle frame for all methods, in order to consider the most challenging outputs and evaluate with the same number of samples despite that 4DiM generates 4x more frames. We take the immediately previous frame to the middle frame in the case where a model generates an even number of frames.
| Method | FID |
| RIFE | 57.68 |
| FILM | 68.88 |
| LDMVFI | 56.28 |
| 4DiM | 52.89 |
| Input views | FILM | RIFE |
| LDMVFI | 4DiM | |
Qualitatively (Fig. 9), we find that 4DiM is competitive against all of these baselines, and does a much better job when the amount of motion is large. Non-generative baselines suffer from severe artifacts in these cases.
Appendix I Ablation for co-training with video data
In our main evaluations, we find that 4DiM trained without video performs slightly better on in-distribution evaluations (e.g., RealEstate10K). This is not too surprising as much of the capacity of the model now needs to be spent to model a more diverse data distribution. Still, to validate our claim that co-training with video is of paramount importance, it is necessary to consider out-of-distribution qualitative comparisons. Here it becomes evident that a model trained without video performs much worse. In particular, such a model does not produce plausible geometry for unknown objects (Fig. 10), whereas the model trained with large-scale video data does not suffer from this issue.
| Input image | Novel view (no video) | Novel view (with video) |
Appendix J Additional 4DiM samples
@@ Appendix I Ablation for co-training with video data
| Input | Output handling reflections from screen, floor, and water. |
| Input | Output handling light, fog, rain etc. |